OpenStudy (anonymous):

A number cube with faces labeled 1–6 is rolled 72 times. About how many times would the result be 4? A. 4 B. 6 C. 8 D. 12 My answer is D

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

the probability of getting a 4 on 1 throw is what do you think Remember there are 6 numbers on the cube

OpenStudy (anonymous):

72 / 6 = 12

OpenStudy (welshfella):

yesd:- prob 1 throw = 1/6 so expected number = 4 on 72 throws = 12

OpenStudy (anonymous):

So it was D like I said?